OpenStudy (anonymous):

2. Simplify each expression. a. (3xy^3)(2x^3y) b. (x^2+2x+3)(x^2-2x-2)

OpenStudy (anonymous):

add the exponents on the like terms for example \(x\times x^3=x^4\)

OpenStudy (anonymous):

These are polynomials.

OpenStudy (anonymous):

I know, I'm horrible with them.

OpenStudy (anonymous):

wanna do a first?

OpenStudy (anonymous):

yes please, I reaally need help, I have a quiz tomorrow.

OpenStudy (anonymous):

then it would be best if you do this yourself i will show you the steps for \[ (3xy^3)(2x^3y)\]what is \(3\times 2\) ?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

ok next what is \(x\times x^3\) ?

OpenStudy (anonymous):

x^4

OpenStudy (anonymous):

I was about to ask her the same. Seems like you have it

OpenStudy (anonymous):

right and finally what is \(y^3\times y\) ?

OpenStudy (anonymous):

y^4

OpenStudy (anonymous):

And now your final answer...

OpenStudy (anonymous):

ok good then put it all together (write them next to each other) and what you do get as a "final answer"

OpenStudy (anonymous):

6+x^4+y^4 ?

OpenStudy (anonymous):

oh no dear it is times, not plus you are not adding

OpenStudy (anonymous):

Careful, you do not want to add them.

OpenStudy (anonymous):

Satellite, you beat me to it.

OpenStudy (anonymous):

ooh so replace the plus with the multiplication symbol.

OpenStudy (anonymous):

or replace them by nothing

OpenStudy (anonymous):

You don't even need a multiplication symbol. Terms that are together are multiplied.

OpenStudy (anonymous):

instead of writing \[6\times x^4\times y^4\] you can just write \[6x^4y^4\]

OpenStudy (anonymous):

oh! okay thanks for help on the first one guys c:

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

Do you know what you have to do with the second one

OpenStudy (anonymous):

kind of haha

OpenStudy (anonymous):

I'll start you out here...

OpenStudy (anonymous):

 Do you understand what the lines mean